A triangular lamina in the Ty-plane such tha
its vertices are (0,0), (0,1) and (1,0). Suppose

that the density function of the lamina is
y) = 63ry g
defined by p(x.
centimetre. Use the information given to
answer the following questions (2 decimal
places):

The total mass of the lamina is
grams.
The moment about the x -axis of the lamina

The Moment about the y -axis of the lamina

The centre of gravity of the lamina ( x, y) is (
1
Expert's answer
2020-05-26T20:04:17-0400

Area of the triangle having vertices (0,0), (1,0) and (0,1) is "\frac{1}{2}" square unit.

The equation of hypotenuse of triangle is "y = -x+1" .

Hence the region D is the triangle enclosed by the lines :

"x =0, y = 0, x+y = 1" .

a) The total mass of the lamina is "m = \int\int_D \rho (x,y) dA".

Given "\rho(x,y) = 63 y".

So, "m = \int_0^1 \int_0^{-x+1} 63 y dydx = \int_0^1 [63 \frac{y^2}{2}]_0^{-x+1} dx = [\frac{63}{2} \frac{(1-x)^3}{-3}]_0^1 = \frac{63}{6}".


b) The center of mass "(\bar{x},\bar{y})" is given by

"\bar{x} = \frac{1}{m} \int\int_D x\rho(x,y)dA , \bar{y} = \frac{1}{m} \int\int_D y\rho(x,y)dA".

Now,

"\int\int_D x\rho(x,y)dA = \int_0^1 \int_0^{-x+1} 63x y dydx = \int_0^1 [63x \frac{y^2}{2}]_0^{-x+1} dx \\
= \frac{63}{2} \int_0^1 x(1-x)^2 dx = \frac{63}{2} \int_0^1 (x^3-2x^2 +x) dx = \frac{63}{2} [\frac{x^4}{4} - \frac{2x^3}{3}+\frac{x^2}{2}]_0^1 = \frac{63}{24}"

and

"\int\int_D y\rho(x,y)dA = \int_0^1 \int_0^{-x+1} 63 y^2 dydx = \int_0^1 [63 \frac{y^3}{3}]_0^{-x+1} dx \\
= \frac{63}{3} \int_0^1 (1-x)^3 dx = \frac{63}{3} [\frac{(1-x)^4}{(-4)}]_0^1 = \frac{63}{12}"

So, "\bar{x} = \frac{6}{63} \times\frac{63}{24} = \frac{1}{4}, \ and \ \bar{y} = \frac{6}{63} \times\frac{63}{12} = \frac{1}{2}" .